the difference of roots of equation px^2+qx+r=0 is

Grade:12

1 Answers

rajat Bansal
9 years ago
suppose roots are ​​ A and B
So we  know the sum of the roots is A +B =-q/p
and multiplication of roots is A*B= r/p
now
Difference between roots
  1. A−B
  2. A−B = ( (A−B)^2 )^1/2
  3. A−B = ( A^2 +B^2 −2*A*B)^1/2
  4. A−B = ( A^2 +B^2 +2*A*B −4*A*B)^1/2                ( ADD 2AB and minus 2AB_)
  5. A−B = ( ( A+B)^2 −4*A*B)^1/2
  6. A−B = ( (-q/p)^2 -4*r/p)^1/2